# Variational Quantum Algorithm Landscape Reconstruction by Low-Rank Tensor Completion

### v1

@@ Abstract
Variational quantum algorithms (VQAs) are a broad class of algorithms with many applications in science and industry. Applying a VQA to a problem involves optimizing a parameterized quantum circuit by maximizing or minimizing a cost function. A particular challenge associated with VQAs is understanding the properties of associated cost functions. Having the landscapes of VQA cost functions can greatly assist in developing and testing new variational quantum algorithms, but they are extremely expensive to compute. Reconstructing the landscape of a VQA using existing techniques requires a large number of cost function evaluations especially when the dimension or the resolution of the landscape is high. To address this challenge, we propose a low-rank tensor-completion-based approach for local landscape reconstruction. By leveraging compact low-rank representations of tensors, our technique can overcome the curse of dimensionality and handle high-resolution landscapes. We demonstrate the power of landscapes in VQA development by showcasing practical applications of analyzing penalty terms for constrained optimization problems and examining the probability landscapes of certain basis states.
@@ II-B Tensor Networks
Tensor networks are a powerful computational tool widely used in Mathematics, Physics, Chemistry, and Computer Science [26, 27, 28, 29, 30, 31]. Below, we provide a brief review of tensor networks, focusing on the concepts used in this paper.
@@
A tensor network is a generalized graph (with dangling edges) of tensors connected by shared indices. In particular, the 1D chain tensor networks, known as Tensor Train (TT) [32] in Mathematics, or Matrix Product States (MPS) [33] in Physics and Chemistry, are very successful in various applications. An $d$-site MPS consists of 2 order-2 tensors $\mathbb{M}^{(1)},\mathbb{M}^{(d)}$ and $d-2$ order-3 tensors $\mathbb{M}^{(2)},\cdots\mathbb{M}^{(d-1)}$, such that the full contraction
@@
gives the dense tensor it represents. The “open” indices $i_{1},\cdots,i_{d}$ that are uncontracted are called physical indices, which usually correspond to basis states of a quantum system, discrete variables, or parameters. The interconnected indices $r_{1},r_{2},\cdots,r_{d-1}$ are called virtual indices or bond indices and their maximum value $R_{1},R_{2},\cdots,R_{d-1}$ are called TT ranks or bond dimensions.
Figure 1: Overview of the proposed landscape reconstruction method by low-rank tensor completion. A small number of quantum circuit outputs are used to perform tensor completion on low-rank representation of the landscape.
@@ III Landscape Reconstruction by Low-rank Tensor Completion
Given a $d$-dimensional VQA parameter $\bm{\theta}=\left(\theta_{1},\ldots,\theta_{d}\right)$ with indices $\left(i_{1},\ldots,i_{d}\right)$, $f(\bm{\theta})$ can be approximated by Eq. (2). The space complexity of a TT representation is $\mathcal{O}(dR^{2}N)$, reducing the memory requirements from exponential to linear with respect to $d$ and $N$. The compact TT representation accurately expresses the landscape as long as it is low rank. We show evidence that the landscape is low-rank in Section III-B. Here, we use the TT format, but other tensor network and tensor decomposition formats are generally applicable.
@@ III-B Evidence that Local VQA Landscapes are Low-rank
To verify that the local VQA landscapes are low-rank, we obtain the landscapes of various configurations in the dense tensor format by numerical simulation. With the dense landscapes, we perform sequential reshaping and singular value decompositions (SVD) along each dimension of the dense tensor to transform it into the TT format. This process is known as the TT decomposition, where the approximation error can be upper bounded by the truncation threshold of the singular values during SVDs. The number of remaining singular values after truncation is the aforementioned bond dimension or TT rank.
@@ III-C Evaluation
We implement the TT representation of landscapes and the tensor completion reconstruction workflow as part of the VQA helper package OSCAR (https://github.com/QUEST-UWMadison/OSCAR). We employ the teneva package [39, 35] for the tensor completion functionalities and provide a unified interface with the compressed-sensing-based reconstruction method in [11].
@@ V Conclusion
We highlight a few directions for future work. From the reconstruction algorithm perspective, it is worth investigating the properties of specific VQAs to further reduce the sampling cost of tensor completion, such as better strategies of the initial sampling and more customized tensor network structure for functional approximation of quantum algorithms. From the application perspective, it would be interesting to apply the landscape reconstruction techniques to study the impact of quantum noise and to perform error mitigation.

### v2

@@ Abstract
Variational quantum algorithms (VQAs) are a broad class of algorithms with many applications in science and industry. Applying a VQA to a problem involves optimizing a parameterized quantum circuit by maximizing or minimizing a cost function. A particular challenge associated with VQAs is understanding the properties of associated cost functions. Having the landscapes of VQA cost functions can greatly assist in developing and testing new variational quantum algorithms, but they are extremely expensive to compute. Reconstructing the landscape of a VQA using existing techniques requires a large number of cost function evaluations, especially when the dimension or the resolution of the landscape is high. To address this challenge, we propose a low-rank tensor-completion-based approach for local landscape reconstruction. By leveraging compact low-rank representations of tensors, our technique can overcome the curse of dimensionality and handle high-resolution landscapes. We demonstrate the power of landscapes in VQA development by showcasing practical applications of analyzing penalty terms for constrained optimization problems and examining the probability landscapes of certain basis states.
@@ II-B Tensor Networks
A tensor network [26, 27, 28, 29, 30, 31] is a generalized graph (with dangling edges) of tensors connected by shared indices. In particular, the 1D chain tensor networks, known as Tensor Train (TT) [32] in Mathematics, or Matrix Product States (MPS) [33] in Physics and Chemistry, are very successful in various applications. A $d$-site MPS consists of 2 order-2 tensors $\mathbb{M}^{(1)},\mathbb{M}^{(d)}$ and $d-2$ order-3 tensors $\mathbb{M}^{(2)},\cdots\mathbb{M}^{(d-1)}$, such that the full contraction
@@
gives the dense tensor it represents. The bond dimensions $R_{1},R_{2},\cdots,R_{d-1}$ of the interconnected indices $r_{1},r_{2},\cdots,r_{d-1}$ are referred to as ranks.
Figure 1: Overview of the proposed landscape reconstruction method by low-rank tensor completion. A small number of quantum circuit outputs are used to perform tensor completion on the low-rank representation of the landscape.
@@ III Landscape Reconstruction by Low-rank Tensor Completion
Given a $d$-dimensional VQA parameter $\bm{\theta}=\left(\theta_{1},\ldots,\theta_{d}\right)$ with indices $\left(i_{1},\ldots,i_{d}\right)$, $f(\bm{\theta})$ can be approximated by Eq. (2). The space complexity of a TT representation is $R_{1}N_{1}+R_{d-1}N_{d}+\sum_{i=2}^{d-1}R_{i-1}R_{i}N_{i}$, reducing the memory requirements from exponential to linear with respect to $d$ and $N_{i}$. The compact TT representation accurately expresses the landscape as long as it is low rank. We show evidence that the landscape is low-rank in Section III-B. Here, we use the TT format, but other tensor network and tensor decomposition formats are generally applicable.
@@ III-B Evidence that Local VQA Landscapes are Low-rank
To verify that the local VQA landscapes are low-rank, we obtain the landscapes of various configurations in the dense tensor format by numerical simulation. With the dense landscapes, we perform sequential reshaping and singular value decompositions (SVD) along each dimension of the dense tensor to transform it into the TT format. This process is known as the TT decomposition, where the approximation error can be upper bounded by the truncation threshold of the singular values during SVDs. The number of remaining singular values after truncation is the aforementioned rank of the TT.
@@ III-C Evaluation
We implement the TT representation of landscapes and the tensor completion reconstruction workflow as part of the VQA helper package OSCAR (https://github.com/QUEST-UWMadison/OSCAR). We employ the teneva package [39, 35] for the tensor completion functionalities and provide a unified interface with the compressed-sensing-based reconstruction method in [11]. Note that due to the discrepancy in the applicable scope of the two methods, it is not possible to directly compare their performance.
@@ V Conclusion
We highlight a few directions for future work. From the reconstruction algorithm perspective, it is worth investigating the properties of specific VQAs to further reduce the sampling cost of tensor completion, such as better strategies of the initial sampling and more customized tensor network structures for functional approximation of quantum algorithms. From the application perspective, it would be interesting to apply the landscape reconstruction techniques to study the impact of quantum noise and to perform error mitigation.
Disclaimer
This paper was prepared for informational purposes with contributions from the Global Technology Applied Research center of JPMorganChase. This paper is not a product of the Research Department of JPMorganChase or its affiliates. Neither JPMorganChase nor any of its affiliates makes any explicit or implied representation or warranty and none of them accept any liability in connection with this position paper, including, without limitation, with respect to the completeness, accuracy, or reliability of the information contained herein and the potential legal, compliance, tax, or accounting effects thereof. This document is not intended as investment research or investment advice, or as a recommendation, offer, or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction.